Assume that Smith Corporation will need to purchase 200,000 British pounds in 90 days. A call option exists on British pounds with an exercise price of $1.68, a 90-day expiration date, and a premium of $.03. A put option exists on British pounds, with an exercise price of $1.69, a 90-day expiration date, and a premium of $.03. Smith Corporation plans to purchase options to cover its future payables. It will exercise the option in 90 days (if at all). It expects the spot rate of the pound to be $1.76 in 90 days. Determine the amount of dollars it will pay for the payables, including the amount paid for the option premium.
Answer #1

As we are expecting the exchange rate to go high we should purchase call option at an exercise price of 1.68

Premium is 0.03

So cost of 1 pound is 1.68+0.03 = 1.71

So amount paid is

= 200000(1.68+0.03) = 342000